Last year, Singer A performed five more than two times as many shows during concert tours than Team B. The number of Singer A concert tour shows exceeded the number of shows by Team B by 23. How many concert shows did each perform?
Answer by ikleyn(52943)

.

Last year . . . 

    A = 2B + 5     (1)


The number . . . 

    A = B + 23.    (2)


Equations (1) and (2) have the same left sides --- hence, their right sides are equal :


    2B + 5 = B + 23.


It implies

    2B - B = 23 - 5

     B     = 18.


ANSWER.  B = 18,  A = B + 23 = 18+23 = 41.

Solved.